At the .05 significance level, does the data below show significant correlation? 

xy
523.6
617.08
720.06
829.84
930.02
1021.1
1120.28
1225.76
1327.24
1431.32
1530.5
1621.38
  • Yes, significant correlation 

  • No
Answer #1
ΣX ΣY Σ(x-x̅)² Σ(y-ȳ)² Σ(x-x̅)(y-ȳ)
total sum 126.00 298.18 143.00 263.82 74.29
mean 10.50 24.85 SSxx SSyy SSxy

correlation coefficient ,    r = SSxy/√(SSx.SSy) =   0.38248

n=   12  
alpha,α =    0.05  
correlation , r=   0.3825  
t-test statistic = r*√(n-2)/√(1-r²) =        1.3090
DF=n-2 =   10  
p-value =    0.2198  
Decison:   P value > α, So, Do not reject Ho  

answer: NO